# Enhancing Agent Learning through World Dynamics Modeling

### v1

@@ Abstract
While large language models (LLMs) have been increasingly deployed across tasks in language understanding and interactive decision-making, their impressive performance is largely due to the comprehensive and in-depth domain knowledge embedded within them. However, the extent of this knowledge can vary across different domains. Existing methods often assume that LLMs already possess such comprehensive and in-depth knowledge of their environment, overlooking potential gaps in their understanding of actual world dynamics. To address this gap, we introduce Discover, Verify, and Evolve (DiVE), a framework that discovers world dynamics from a small number of demonstrations, verifies the correctness of these dynamics, and evolves new, advanced dynamics tailored to the current situation. Through extensive evaluations, we analyze the impact of each component on performance and compare the automatically generated dynamics from DiVE with human-annotated world dynamics. Our results demonstrate that LLMs guided by DiVE can make better decisions, achieving rewards comparable to human players in the Crafter environment.
@@
Zhiyuan Sun1thanks:   Equal contribution., Haochen Shi [^1] footnotemark: 1, Marc-Alexandre Côté2, Glen Berseth1, Xingdi Yuan2, Bang Liu1 1 Université de Montréal & Mila, Montréal, Canada 2 Microsoft Research, Montréal, Canada
@@ 1 Introduction
By absorbing internet-scale knowledge autoregressively, Large Language Models (LLMs) develop a nuanced understanding of world dynamics Achiam et al. (2023); Team et al. (2023); Brown et al. (2020); Touvron et al. (2023). This understanding enables them to perform well on tasks like question answering Yang et al. (2018); Clark et al. (2019); Lin et al. (2021), planning Song et al. (2023); Brown et al. (2020); Huang et al. (2022), and reasoning over commonsense knowledge Yao et al. (2022, 2024); Besta et al. (2024). However, such understanding requires the training dataset to be diverse, precise, and in-depth enough to cover the domain information needed; otherwise, this may result in a knowledge gap between LLMs and the target domain, as illustrated in Figure 1.
As LLMs tend to follow the most common patterns within the dataset Gunasekar et al. (2023), this mandates that the necessary domain-specific information not only appears on the internet but also appears frequently. For downstream tasks that are either newly established or less popular, such a requirement is often infeasible and leads to suboptimal decision-making for LLMs.
Figure 1: The knowledge gap between LLMs and the domain. Although LLMs are trained on internet-level knowledge, they may still lack the necessary domain-specific knowledge.
To make optimal decisions, LLMs require high-quality, domain-specific information, which is often impractical to source and verify within the vast expanse of internet data Dziri et al. (2022a). Evaluating such information necessitates locating it within the dataset and verifying it against ground-truth domain knowledge, a process that is both labor-intensive and error-prone. Consequently, ensuring LLMs possess the necessary comprehensive and accurate knowledge for decision-making remains a formidable challenge.
@@
Even if LLMs have a general understanding of a downstream domain, achieving optimal decision-making requires in-depth knowledge specific to each state. By providing LLMs with such tailored and in-depth information, their understanding of both the environment and the current state can be further improved. For example, mastering the game of Go involves not only understanding the basic rules but also strategies tailored to specific board states. Moreover, this specialized knowledge can vary not only across domains but also within different states of the same domain, making it impractical to collect in an offline manner.
@@
To address these challenges, we propose Discover, Verify, and Evolve (DiVE). Leveraging the concept of World Model Ha and Schmidhuber (2018); Hafner et al. (2020, 2023), DiVE not only discovers and verifies the primitive world dynamics from the demonstrations but also evolves state-specific knowledge for the downstream domain. By providing LLMs with all these dynamics and knowledge, DiVE can bridge the understanding gap between LLMs and the downstream domain, thereby enhancing LLMs’ decision-making abilities.
DiVE comprises three distinct components: i) The Discoverer: This component iteratively uncovers the dynamics of the environment based on the provided demonstrations. It begins by discovering the primitive dynamics of actions and objects, progresses to subtask steps, and finally determines the topological order of subgoals. This curriculum learning method ensures both learning efficiency and a comprehensive understanding of the environment’s dynamics. ii) The Verifier: This component filters out inaccuracies caused by the tendency of LLMs to hallucinate, ensuring that only precise and reliable information is retained. This step is crucial for maintaining the integrity of the knowledge base. iii) The Evolver: This component reasons in-depth, state-appropriate dynamics, such as strategic plays or contextual decisions. It ensures that the agent is equipped with actionable knowledge tailored to the current state, enhancing real-time decision-making.
@@
Within the Crafter environment Hafner (2021), DiVE learns comprehensive and accurate dynamics from demonstrations and guides the agent through the decision-making process by developing in-depth knowledge. This guidance enables the agent to outperform strong baselines, including the best reported systems, and achieve rewards comparable to those of human players. We also provide both quantitative and qualitative analyses that help understanding DiVE’s behavior, demonstrating the effectiveness of the proposed framework.
In summary, our primary contribution is a general framework that autonomously learns world dynamics from demonstrations and guides LLMs in decision-making by evolving contextual knowledge. This approach not only enhances the capabilities of LLMs in specific domains but also provides a versatile foundation for creating agents capable of mastering complex tasks across various environments through continuous learning and refinement.
2 Problem Formulation
@@
Figure 2: Overall pipeline of DiVE. Left: Learning basic game dynamics from offline demonstrations (Section 3.1). We want to highlight the incorrect game dynamics being identified by the Verifier (labeled by $\times$), they are evidence of the LLM hallucinates false facts perhaps because of memorizing Minecraft data. Right: Learning situational strategies from online interactions (Section 3.2). For simplicity, we omit the verbalization process in the right figure.
Environment with Hierarchical Achievements
Each environment with hierarchical achievements, $C_{i}\in\mathbbm{C}$, operates interactively. At every time step $t$, an agent observes $o_{t}\in\mathcal{O}$ and is expected to respond by selecting an action $a_{t}\in\mathcal{A}$. Subsequently, the environment returns a new observation $o_{t+1}$ of the resulting state. Within each $C_{i}$, the achievement graph $\mathcal{G}=(V,E)$ is structured as a directed acyclic graph (DAG). In this graph, each vertex $v\in V$ denotes an achievement and each edge $(u,v)\in E$ indicates that the achievement $v$ has a dependency on the achievement $u$. To unlock achievement $v$, one must first unlock all its prerequisite ancestors. The main objective is to maximize the number of achievements unlocked by satisfying these dependency conditions.
The Crafter Environment
We primarily employ the Crafter environment as a benchmark to evaluate an agent’s performance in a setting characterized by hierarchical achievements. The Crafter environment  Hafner (2021) is a long-horizon environment typically necessitating hundreds of steps to successfully collect a diamond. Within this environment, the agent must not only determine the topological order of the technology tree as illustrated in Appendix E, but also survive by overcoming dangerous creatures and managing diminishing resources.
Challenge: The Knowledge Gap
In this work, we assume a setting where a pre-trained LLM-based agent is used to solve tasks in the Crafter environment. Conceptually, we use $\mathcal{K}_{\text{data}}$ to denote the set of knowledge being embedded in the LLM through its training process (including both pre-training and fine-tuning). We use $\mathcal{K}_{\text{target}}$ to denote the universal set of knowledge required to solve the target domain (i.e., Crafter). For the LLM to be useful, we hope $\mathcal{K}_{\text{rel}}$, the subset of knowledge in $\mathcal{K}_{\text{data}}$ that is relevant to $\mathcal{K}_{\text{target}}$ to be as large as possible. We also hope that $\mathcal{K}_{\text{rel}}$ contains more accurate knowledge $\mathcal{K}_{+}$ than inaccurate knowledge $\mathcal{K}_{-}$, where $\mathcal{K}_{\text{rel}}=\mathcal{K}_{+}\cup\mathcal{K}_{-}$. Therefore, we define three desirable properties to help conceptualize our system design:
@@
  • Recall $R=\frac{|\mathcal{K}_{+}|}{|\mathcal{K}_{\text{target}}|}$ measures the extent of knowledge required to solve tasks in the target domain is being covered by the LLM’s knowledge. A low recall typically suggests a large distribution gap between LLM’s training data and the target domain.
@@
  • Precision $P=\frac{|\mathcal{K}_{+}|}{|\mathcal{K}_{\text{rel}}|}$ measures how accurate the LLM’s knowledge is when grounded in the target domain. An example of inaccurate game dynamics is shown in Figure 2 (left), where the LLM hallucinates that defeating skeleton will drop items, which is not true in Crafter. This may because the LLM has memorized some Minecraft-specific data.
@@
  • Depth $D$ measures the abstraction levels of the knowledge representations. Knowledge can be represented at different abstraction levels, ranging from elementary level game rules to high level strategies.
We want to acknowledge that it is nontrivial to directly quantify the exact knowledge overlap between an LLM and a dataset. We adopt the above concepts from Information Retrieval and related literature to guide our system design. In Section 3, we show mathematically how our system improve on all the three dimensions; in Section 4, we design silver standard on quantifying the knowledge overlap, which provides us evidence that our system can indeed learn useful game dynamics and to some extent fills the knowledge gap.
3 Method
@@
In an ideal world, one could alleviate the aforementioned knowledge gap by fine-tuning the LLM to adapt to the target domain. However, this approach often show less practical due to its reliance on abundant annotated data and significant computational overhead Hu et al. (2021); Zheng et al. (2024); Carta et al. (2023); Ouyang et al. (2022). Our framework, DiVE, is designed to bridge the knowledge gap taking into consideration all the three desirable properties (i.e., recall, precision, and depth), but without the need of collecting extensive data from the target domain. It is a prompt-based method that learns world dynamics $\mathcal{W}$ from the environment.
@@
As illustrated in Figure 2, DiVE is initially bootstrapped from a handful of human demonstration trajectories $H=\{\tau_{i}\}_{i=0}^{N}=\{\{o_{t}^{i},a_{t}^{i},r_{t}^{i}\}_{t=0}^{|\tau_{i}|}\}_{i=0}^{N}$. We verbalize an observation $o_{t}$ into the language space as $\tilde{o}_{t}$ using a Verbalizer (i.e., $\tilde{o}_{t}=\text{Verbalizer}(o_{t})$), resulting in transformed trajectories $\tilde{\tau}_{i}=\{\tilde{o}_{t}^{i},a_{t}^{i},r_{t}^{i}\}_{t=0}^{|\tau_{i}|}$. Next, the Discoverer distills a set of world dynamic candidates $\tilde{\mathcal{W}}=\{\tilde{W}_{+},\tilde{W}_{-}\}$ from human demonstrations $H$, where $\tilde{W}_{+}$ and $\tilde{W}_{-}$ represent the correct and inaccurate world dynamic sets, respectively.
@@
Empirically, we find that it is often inevitable to have $\tilde{W}_{-}$ in $\tilde{\mathcal{W}}$, either due to the backbone LLM’s imperfection in discovering meaningful knowledge from trajectory data, or its tendency of hallucination. Therefore, we use a Verifier to filter out potential invalid and conflicting world dynamic candidates within $\tilde{\mathcal{W}}$, we denote the remaining valid dynamics as $\mathcal{W}$. Next, we use an Evolver, which is designed to derive advanced game-playing strategies $\mathcal{I}$ based on the world dynamics $\mathcal{W}$ and the language description of observation $\tilde{o}_{t}$.
@@
The final decision-making process on primitive actions $a_{t}\in\mathcal{A}$ is hierarchically decomposed as planning tasks on sub-goals $SG$, sub-tasks $ST$, and actions $\mathcal{A}$. The planning procedure is further guided by both $\mathcal{W}$ and $\mathcal{I}$. In cases where $\mathcal{W}\neq\varnothing$, $R$, $P$, and $D$ are guaranteed to increase as formulated below:
@@
The DiVE framework can be split into two stages: an offline dynamics learning stage and an online strategy learning stage.
@@ 3.1 Offline Dynamics Learning
The offline dynamics learning procedure aims at bridging the gap between LLM’s understanding and the basic rules of the Crafter world by learning the world dynamics $\mathcal{W}$ as a prior for further decision-making. Rather than extracting world dynamics from human-authored game manuals or handbooks  Wu et al. (2024b, a), which can be difficult to obtain in many real-world scenarios, we explore learning world dynamics $\mathcal{W}$ from experiences $H$, which is arguably more accessible.
@@
Given the varying complexities in learning the dynamics of different elements in Crafter, we adopt curriculum learning Bengio et al. (2009). We propose a sequential learning approach progressing from simple to more challenging dynamics, facilitating effective learning. Specifically, we propose a method to learn the dynamics of each element within the task decomposition hierarchy, denoted as $TD=\{\mathcal{A}+(O),ST,SG\}$, where $O$ is the set of objects in Crafter. Our approach starts with elements of lower abstraction, such as actions $a\in\mathcal{A}$, and progresses to higher abstraction elements, like sub-goals $sg_{i}\in SG$. The sub-goals sequence $SG=[sg_{1},sg_{2},\ldots]$ is an ordered sequence to unlock achievements in the achievement graph $\mathcal{G}=(V,E)$, with $SG=\text{TopologicalSort}(\mathcal{G})$ and each $sg_{i}$ being a vertex in $V$. We use the Discoverer to extract this order from human demonstrations $H$. The sub-tasks $ST$ are defined as the nodes in the achievement graph $G$, i.e., $ST=V$. Achieving a sub-goal $sg_{i}$ may require completing several sub-tasks multiple times. This approach ensures a logical progression through the tasks, enhancing the learning and application of Crafter’s dynamics.
@@
The Discoverer is designed to identify dynamic candidates $\tilde{\mathcal{W}}$ relating to the elements within the task decomposition hierarchy $TD$. A single dynamics discovering step for an element $E\in TD$ involves following three main steps:
@@
    Construction of a Semantic Experience Bank ($B^{E}$): For each element $E$, a semantic experience bank $B^{E}$ is constructed from human demonstrations $H$. This bank stores experiences that are transformed from $H$ into a suitable granularity for analyzing dynamics related to $E$. The transformation involves chunking and summarizing the verbalized human demonstrations $\tilde{H}$ to capture essential semantic details.
@@
The action-level semantic experience bank, denoted as $B^{\mathcal{A}}$, stores transition tuples derived from verbalized human demonstrations $\tilde{H}$, specifically $B^{\mathcal{A}}=\{\{\tilde{o}t,a_{t},\tilde{o}{t+1}\}_{i}\}_{i=1}^{|B^{\mathcal{A}}|}$. Similarly, the object-level semantic experience bank $B^{O}$ collects individual observations related to specific object interactions from $\tilde{H}$, stored as $B^{O}=\{\tilde{o}i\}_{i=1}^{|B^{O}|}$. The sub-task-level semantic experience bank $B^{ST}$ aggregates trajectory segments that represent the completion of sub-tasks, formatted as $B^{ST}=\{\{\tilde{o}_{t},\ldots,a_{tst},\tilde{o}_{tst+1}\}_{i}\}_{i=1}^{|B^{ST}|}$, where $tst$ denotes the time step at which a sub-task $st\in ST$ is completed. For sub-goals, the sub-goal-level experience bank $B^{SG}$ records sequences of sub-tasks that lead to the completion of sub-goals, expressed as $B^{SG}=\{\{st_{t},\ldots,st_{tsg}\}_{i}\}_{i=1}^{|B^{SG}|}$, where $tsg$ is the time step at which the sub-goal $sg$ is achieved.
For action-level dynamics learning, the relevant experiences $B_{a}^{\mathcal{A}}$ are compiled by randomly sampling transition tuples from $B^{\mathcal{A}}$ where the action $a$ is successfully executed. A similar approach is applied to the dynamic learning for other elements within the task decomposition hierarchy $TD$.
For action-level dynamics discovery, we identify the prerequisites and outcomes of each action (e.g., The action MakeWoodPickaxe requires 1 wood). For object-level dynamics, we concentrate on the co-occurrence relationships between objects and their occurrences over time. The attribute set for a sub-task typically includes the general steps required to complete the task, along with its prerequisites, outcomes, and termination conditions. The only attribute of interest for a sub-goal is its correct position within the sub-goal sequence $SG$.
@@
Dynamic discovery processes are susceptible to noise arising from several sources, such as confounders, hallucinations by the LLM, and imperfections in the LLM’s ability to derive meaningful knowledge from trajectory data. To mitigate these issues, we introduce a dynamic verifier designed to filter out all noisy dynamic candidates from $\tilde{\mathcal{W}}$. For each dynamic candidate $w\in\tilde{\mathcal{W}}$, the Verifier initiates the process by sampling a subset of relevant semantic experiences $B_{e}^{E}$ from the corresponding semantic experience bank $B_{E}$. In this context, $w$ represents a dynamic candidate related to one attribute of the instance $e\in E$, where $E\in\{\mathcal{A}+(O),ST,SG\}$ is an element of the task decomposition hierarchy $TD$. The verification of $w$ proceeds as follows: $w$ is identified as inaccurate and consequently filtered out if it does not consistently hold across all experiences within $B_{e}^{E}$ or if it conflicts with any other established dynamics. The dynamics that withstand this verification process are classified as verified dynamics, denoted as $\mathcal{W}$.

### v2

@@ Abstract
Large language models (LLMs) have been increasingly applied to tasks in language understanding and interactive decision-making, with their impressive performance largely attributed to the extensive domain knowledge embedded within them. However, the depth and breadth of this knowledge can vary across domains. Many existing approaches assume that LLMs possess a comprehensive understanding of their environment, often overlooking potential gaps in their grasp of actual world dynamics. To address this, we introduce Discover, Verify, and Evolve (DiVE), a framework that discovers world dynamics from a small number of demonstrations, verifies the accuracy of these dynamics, and evolves new, advanced dynamics tailored to the current situation. Through extensive evaluations, we assess the impact of each component on performance and compare the dynamics generated by DiVE to human-annotated dynamics. Our results show that LLMs guided by DiVE make more informed decisions, achieving rewards comparable to human players in the Crafter environment and surpassing methods that require prior task-specific training in the MiniHack environment111  The code is available at https://github.com/ZhiyuuanS/DiVE.
@@
Zhiyuan Sun1thanks:   Equal contribution. thanks:   zhiyuan.sun@umontreal.ca, Haochen Shi [^1] footnotemark: 1 Marc-Alexandre Côté2, Glen Berseth1,3, Xingdi Yuan2thanks:   Equal advising., Bang Liu1, [^3] footnotemark: 3 1 Université de Montréal & Mila, Montréal, Canada 2 Microsoft Research, Montréal, Canada 3 Canada CIFAR AI Chair
@@ 1 Introduction
By absorbing internet-scale knowledge autoregressively, large language models (LLMs) develop a broad understanding of the world Achiam et al. (2023); Team et al. (2023); Brown et al. (2020); Touvron et al. (2023). This understanding enables them to perform well across a variety of tasks Brown et al. (2020); Huang et al. (2022); Yao et al. (2022, 2024). However, achieving this level of comprehension requires a training dataset that is diverse, precise, and in-depth to cover essential domain information. Otherwise, a knowledge gap may emerge between LLMs and the target domain, as illustrated in Figure 1.
Since LLMs tend to follow the most common patterns in the dataset Gunasekar et al. (2023) and are prone to hallucination, it is crucial that the necessary information not only appears frequently on the internet but is also reliable Dziri et al. (2022). For tasks and environments that are newly emerging Hafner (2021); Samvelyan et al. (2021), these requirements are often difficult to meet due to the noisy nature of internet data, leading to potential knowledge gaps.
Figure 1: The knowledge gap between LLMs and downstream domains. Although LLMs have a broad understanding of the world, they may struggle to grasp the complex dynamics of specific downstream domains.
@@
While LLMs may have a general understanding of a domain, optimal decision-making requires in-depth, state-specific knowledge. Providing LLMs with such tailored information enhances their grasp of both the environment and the current state. For example, mastering Go involves not just knowing the rules but also applying strategies for specific board states. This specialized knowledge varies across domains and states, making offline collection impractical.
@@
To address these challenges, we propose Discover, Verify, and Evolve (DiVE). Building on the concept of the World Model Ha and Schmidhuber (2018), DiVE not only discovers and verifies primitive world dynamics from demonstrations but also evolves state-specific knowledge for the downstream domain. By providing LLMs with this comprehensive set of dynamics, DiVE bridges the knowledge gap between LLMs and the downstream domain, thereby enhancing their decision-making abilities.
DiVE consists of three distinct components: The Discoverer: this component iteratively uncovers the environment’s dynamics from demonstrations using a curriculum learning approach. The Verifier: this component eliminates unreliable dynamics caused by LLMs’ tendency to hallucinate. The Evolver: this component reasons through in-depth, state-specific strategies tailored to the current situation based on the learned dynamics.
@@
In the Crafter environment Hafner (2021) and the MiniHack environment Samvelyan et al. (2021), DiVE learns comprehensive and reliable dynamics from demonstrations, guiding the agent’s decision-making process by evolving in-depth strategies. This enables the agent to outperform all baselines, achieving rewards comparable to human players in the Crafter environment and surpassing the performance of methods requiring task-specific training in the MiniHack environment. To gain deeper insights into DiVE’s behavior, we provide both quantitative and qualitative analyses.
In summary, our primary contribution is a framework that learns world dynamics from demonstrations, guiding LLMs in the decision-making process by online evolving contextual strategies. This approach bridges potential knowledge gaps, resulting in a more optimal decision-making process for LLMs.
2 The Knowledge Gap
@@
Figure 2: Overall pipeline of DiVE. Left: Learning basic game dynamics from offline demonstrations (Section 3.1). We want to highlight the incorrect game dynamics being identified by the Verifier (labeled by $\times$), they are evidence of the LLMs hallucinate false facts perhaps because of memorizing Minecraft data. Right: Learning situational strategies from online interactions (Section 3.2). For simplicity, we omit the verbalization process in the right figure.
In this work, we consider a setting where a pre-trained LLM-based agent is employed to solve tasks in downstream domains. Conceptually, we define $\mathcal{K}_{\text{LLM}}$ as the set of knowledge embedded in the LLMs through their training process, and $\mathcal{K}_{\text{target}}$ as the universal set of knowledge relevant to the downstream domains.
To ensure the effectiveness of the LLMs, we aim for $\mathcal{K}_{\text{relevant}}$, the subset of $\mathcal{K}_{\text{LLM}}$ that is relevant to $\mathcal{K}_{\text{target}}$, to cover the broadest possible scope. Additionally, we seek for $\mathcal{K}_{\text{relevant}}$ to contain more reliable knowledge $\mathcal{K}_{+}$ than inaccurate knowledge $\mathcal{K}_{-}$, where $\mathcal{K}_{\text{relevant}}=\mathcal{K}_{+}\cup\mathcal{K}_{-}$. Thus, we define three desirable properties to guide our system design:
@@
  • Recall $R=\frac{|\mathcal{K}_{+}|}{|\mathcal{K}_{\text{target}}|}$ measures the extent to which the knowledge required to solve tasks in the target domain is covered by the LLMs. A low recall typically indicates a significant knowledge gap between the LLMs’ training data and the target domain.
@@
  • Precision $P=\frac{|\mathcal{K}_{+}|}{|\mathcal{K}_{\text{relevant}}|}$ measures the accuracy of the LLMs’ knowledge when applied to the target domain. An example of inaccurate dynamics is shown in Figure 2 (left), where the LLMs hallucinate that defeating a skeleton will drop items, which is not true in Crafter. This may be due to the LLMs memorizing Minecraft-specific data.
@@
  • Depth $D$ measures the abstraction levels of knowledge representations. Knowledge can be represented at varying levels of abstraction, from basic game rules to higher-level strategies.
We recognize that precisely quantifying the knowledge overlap between LLMs and a downstream domain is challenging. In Section 3, we provide a mathematical analysis showing how DiVE improves across all three dimensions. In Section 6, we measure the knowledge overlap to demonstrate that DiVE effectively learns useful game dynamics and, to some extent, bridges the knowledge gap.
3 DiVE: Discover, Verify, and Evolve
@@
In an ideal scenario, one could bridge the knowledge gap by fine-tuning the LLM to adapt to the target domain. However, this approach is often less practical due to its reliance on large amounts of annotated data and significant computational overhead Hu et al. (2021); Zheng et al. (2024); Carta et al. (2023); Ouyang et al. (2022). Our framework, DiVE, is designed to address the knowledge gap while considering all three desirable properties—recall, precision, and depth—without requiring extensive data collection from the target domain. It is a prompt-based method that learns world dynamics $\mathcal{W}$ directly from the environment.
@@
As shown in Figure 2, DiVE is initially bootstrapped using a small set of human demonstration trajectories $H$, each consisting of observations $o_{t}$, actions $a_{t}$, and rewards $r_{t}$ at each timestep $t$. We then transform each observation $o_{t}$ into the language space as $\tilde{o}_{t}$ using a Verbalizer, resulting in trajectories represented by $\tilde{o}_{t}$, $a_{t}$, and $r_{t}$ at each timestep $t$. Subsequently, the Discoverer extracts a set of world dynamic candidates, $\tilde{\mathcal{W}}=\{\tilde{W}_{+},\tilde{W}_{-}\}$, from human demonstrations $H$, where $\tilde{W}_{+}$ and $\tilde{W}_{-}$ represent the correct and inaccurate world dynamic sets, respectively.
@@
Empirically, we find that the inclusion of $\tilde{W}_{-}$ in $\tilde{\mathcal{W}}$ is often unavoidable, either due to LLM’s difficulties in extracting meaningful knowledge from trajectory data or its tendency to hallucinate. To address this, we employ the Verifier to filter out potentially invalid and conflicting world dynamic candidates from $\tilde{\mathcal{W}}$, leaving only the valid dynamics $\mathcal{W}$. Lastly, we use the Evolver, designed to derive advanced dynamics $\mathcal{I}$ tailored to the verbalized observation $\tilde{o}_{t}$ based on the filtered world dynamics $\mathcal{W}$.
@@
The final decision-making process on primitive actions $a_{t}\in\mathcal{A}$ is hierarchically decomposed as planning tasks on sub-goals $SG$, sub-tasks $ST$, and actions $\mathcal{A}$. The planning procedure is further guided by both $\mathcal{W}$ and $\mathcal{I}$. In cases where $\mathcal{W}\neq\varnothing$, Recall, Precision and Depth are guaranteed to increase as formulated below:
@@
The DiVE framework can be divided into two stages: an offline dynamics learning stage and an online strategy learning stage.
@@ 3.1 Offline Dynamics Learning
The offline dynamics learning procedure aims to bridge the knowledge gap between LLMs’ understanding and the basic dynamics of downstream domains by learning the world dynamics $\mathcal{W}$ as a prior for the decision-making process. Instead of relying on human-authored game manuals or handbooks to extract world dynamics, as in Wu et al. (2024b, a), which are not only difficult to obtain in many real-world scenarios but also often lack critical details (as demonstrated in Table 6), we propose learning world dynamics $\mathcal{W}$ directly from experiences $H$, which are more accessible and provide richer information.
@@
Given the varying complexities of learning the dynamics of different elements in downstream domains, we adopt a curriculum learning approach Bengio et al. (2009). Our method follows a sequential learning strategy that progresses from simpler to more complex dynamics, thereby enabling more effective learning. Specifically, we propose a method for learning the dynamics of each element within the task decomposition hierarchy, denoted as $TD=\{\mathcal{A}\cup O,ST,SG\}$, where $O$ represents the set of objects in downstream domains.
Our approach begins with elements of lower abstraction, such as actions $a\in\mathcal{A}$ and objects $o\in O$, and gradually progresses to higher-level elements, such as sub-tasks $st_{i}\in ST$. The sub-tasks $ST$ are represented as nodes in the achievement graph $\mathcal{G}=(V,E)$ within the downstream domains, i.e., $ST=V$. Finally, we transition to the subgoals $sg_{i}\in SG$. The subgoal sequence $SG=[sg_{1},sg_{2},\ldots]$ is an ordered list used to unlock achievements in the achievement graph $\mathcal{G}$, where $SG=\text{TopologicalSort}(\mathcal{G})$, and each $sg_{i}$ corresponds to a vertex in $V$.
We leverage the Discoverer to extract this order from human demonstrations $H$. Achieving a sub-goal $sg_{i}$ may involve completing several sub-tasks multiple times. This approach ensures a logical progression through tasks, thereby enabling a deeper understanding and integration of downstream domain dynamics.
@@
The Discoverer is designed to identify dynamic candidates $\tilde{\mathcal{W}}$ related to elements within the task decomposition hierarchy $TD$. A single dynamics discovery step for an element $E\in TD$ involves the following three main steps:
@@
    Construction of the Semantic Experience Bank: For each element $E$, we construct a semantic experience bank $B^{E}$ using demonstrations $H$. This bank stores experiences that are transformed from $H$ into a suitable granularity for analyzing dynamics related to $E$. The transformation process involves chunking and summarizing the verbalized demonstrations to capture essential semantic details.
@@
The action-level semantic experience bank, denoted as $B^{\mathcal{A}}$, stores transition tuples derived from verbal demonstrations and is represented as: $B^{\mathcal{A}}=\{\{\tilde{o}_{t},a_{t},\tilde{o}_{t+1}\}_{i}\}_{i=1}^{|B^{\mathcal{A}}|}$. Similarly, the object-level semantic experience bank, denoted as $B^{O}$, gathers individual observations contain an specific object and is represented as: $B^{O}=\{\tilde{o}_{i}\}_{i=1}^{|B^{O}|}$. The sub-task-level semantic experience bank, denoted as $B^{ST}$, aggregates trajectory segments representing the completion of sub-tasks and is formatted as: $B^{ST}=\{\{\tilde{o}_{t},\ldots,a_{t_{st}},\tilde{o}_{t_{st}+1}\}_{i}\}_{i=1}^{|B^{ST}|}$, where $t_{st}$ denotes the timestep at which a sub-task $st\in ST$ is completed. For sub-goals, the sub-goal-level experience bank, denoted as $B^{SG}$, records sequences of sub-tasks that culminate in the completion of sub-goals, expressed as: $B^{SG}=\{\{st_{t},\ldots,st_{t_{sg}}\}_{i}\}_{i=1}^{|B^{SG}|}$, where $t_{sg}$ is the timestep at which the sub-goal $sg$ is achieved.
For action-level dynamics learning, the relevant experiences, denoted as $B_{a}^{\mathcal{A}}$, are compiled by randomly sampling transition tuples from $B^{\mathcal{A}}$ where the action $a$ has been successfully executed. A similar approach is employed for dynamics learning of other elements within the task decomposition hierarchy $TD$.
For action-level dynamics discovery, we identify the prerequisites and outcomes of each action (e.g., The action MakeWoodPickaxe requires 1 wood). For object-level dynamics, we focus on co-occurrence relationships between objects and their temporal patterns. The attribute set for a sub-task generally encompasses the necessary steps for completion, as well as its prerequisites, outcomes, and termination conditions. In contrast, the primary attribute of interest for a sub-goal is its correct position within the sub-goal sequence $SG$.
@@
Dynamic discovery processes are vulnerable to noise from various sources, including confounders, hallucinations by the LLMs, and difficulties in the LLMs’ ability to extract meaningful insights from trajectory data. To address these challenges, we introduce a dynamic verifier designed to filter out noisy dynamic candidates from $\tilde{\mathcal{W}}$. For each dynamic candidate $w\in\tilde{\mathcal{W}}$, the verifier begins by sampling a subset of relevant semantic experiences, denoted as $B_{e}^{E}$, from the corresponding semantic experience bank $B_{E}$. Here, $w$ represents a dynamic candidate associated with a specific attribute of the instance $e\in E$, where $E\in\{\mathcal{A}\cup O,ST,SG\}$ corresponds to an element of the task decomposition hierarchy $TD$. The verification of $w$ is conducted as follows: $w$ is deemed inaccurate and filtered out if it does not consistently hold across experiences within $B_{e}^{E}$ or if it conflicts with any established dynamics. The dynamics that pass this verification process are classified as verified dynamics and are denoted as $\mathcal{W}$.